1) What is the sum of all odd numbers upto 100?
2) A field is owned by 5 people.A has 3/7 of it and B has twice as much as C.What fractions of the field have B and C respectively?

Answer by drk(1908)

1) What is the sum of all odd numbers upto 100?
1 = 1
1+3 = 4 = 2^2
1+3+5 = 9 = 3^2
1+3+5+ . . . +97+99 = 50^2 = 2500
---
2) A field is owned by 5 people. A has 3/7 of it and B has twice as much as C. What fractions of the field have B and C respectively?
A = (3/7)F
B = 2C
C = C
we know that B + C + D + E = (5/7)F
each person B, C, D, and E has an average of (5/28)F
and
3C + (10/28)F = (20/28)F
3C = (10/28)F
Without knowing anything about D and E, I would say
10/28.